The Z-score uses multiple corporate income and balance sheet values to measure the financial health of a company. Assess the credit risk (z-score) of a potential borrowing firm with the following financial ratios: X1 = 0.2, X2 = 0, X3 = 0.2, X4 = 0.1, X5 = 2.0
1
Expert's answer
2020-06-15T11:11:53-0400

The model is as follows:

Z= 1.2 (X1) + 1.4 (X2) +3.3(X3) + 0.6 (X4) + 1.0 (X5)


Where,

X1= Net working capital/ Total assets

X2= Retained earnings/ Total assets

X3= EBIT/ Total assets

X4= Market value of equity/ Book value of total liabilities

X5= Sales/ Total sales

Z> 2.9 - zone of financial stability ("green" zone).

1.8 <Z <2.9 — the zone of uncertainty (the “gray” zone).

Z <1.8 - zone of financial risk (“red” zone).


X1 = 0,2, X2 = 0, X3 = 0,2, X4 = 0,1, X5 = 2,0


"Z=1.2\times 0.2+1.4\times 0+3.3\times 0.2+0.6\times 0.1+1.0\times 2=0.24+0+0.66+0.06+2=2.96"


potential borrowing firm is in zone of financial stability ("green" zone).